OpenStudy (gabylovesyou):

Determine the type of boundary line and shading for the graph of the inequality 5x + y >/= -10 Dashed line with shading on the side that includes the origin Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

there's a line under the >, so are we dealing with a dashed or a solid line?

OpenStudy (gabylovesyou):

solid

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so that rules out A and C

jimthompson5910 (jim_thompson5910):

so we're left with B and D

jimthompson5910 (jim_thompson5910):

to find out if the origin is in the shaded/solution region, plug in x = 0 and y = 0. If you get a true statement or inequality, then the origin (0,0) is in the shaded region

OpenStudy (gabylovesyou):

B?

jimthompson5910 (jim_thompson5910):

what do you get when you plug x = 0 and y = 0 into 5x + y >= -10

OpenStudy (gabylovesyou):

5*0 + 0 >= -10

jimthompson5910 (jim_thompson5910):

then simplify the left side

OpenStudy (gabylovesyou):

0

jimthompson5910 (jim_thompson5910):

so 5*0 + 0 >= -10 becomes 0 >= -10

OpenStudy (gabylovesyou):

so its D.... it does include it

jimthompson5910 (jim_thompson5910):

yes because 0 >= -10 is true